OpenStudy (anonymous):

Find the center and radius of the circle given the circle x^2+y^2-2ax=0 is tangent to the y-axis at the origin .

OpenStudy (anonymous):

yes that's the equation it has an A in it \[x^2+y^2-2Ax=0\]

OpenStudy (loser66):

Complete the square for x, we have \(x^2 -2ax + a^2 - a^2 +y^2 =0\\(x-a)^2 + y^2 = a^2\) Hence center is (a, 0) , radius a

OpenStudy (anonymous):

o wow that's wat I did then I taugh no way this must be wrong ...thanks